OpenStudy (anonymous):

Explain why csc 180 degree is underfined

OpenStudy (anonymous):

oh yes okay now \(cosec \theta = \frac{H}{P}\)

OpenStudy (anonymous):

Now when \(\theta \rightarrow 180\)

OpenStudy (anonymous):

Since sin 180 = 0 csc 180 = 1/ sin 180 = 1/0 1/0 is undefined

OpenStudy (anonymous):

You can think now why it's not defined

OpenStudy (anonymous):

ah Thanks : )